# Cascade switching current detectors based on arrays of Josephson junctions

### v1

@@ Abstract
Cascade multiplication is a common technique to enhance the sensitivity of photon detectors. In this study, we demonstrate novel cascade-amplified superconducting detectors utilizing arrays of Josephson junctions. The mutual coupling between junctions induces avalanche-like switching of multiple junctions upon photon absorption, leading to cascade amplification of the readout voltage. We present two prototypes featuring either low-$T_{c}$ linear $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Both devices exhibit clear antenna effects indicating good impedance matching and absorption efficiency. The combination of high absorption efficiency and large cascade amplification has the potential to produce broadband THz sensors with sensitivity exceeding $10^{13}\ \mathrm{V/W}$.
@@ Introduction
THz detectors find diverse applications, including security, environmental monitoring, medical imaging, chemical analysis, and fundamental research [1, 2, 3, 4]. Cooling to cryogenic temperatures improves detector sensitivity by reducing thermal noise and background radiation. Ultra-sensitive cryogenic detectors based on semiconductors [5], graphene [6], and superconductors [7, 8, 9, 10, 11, 12, 13] have been demonstrated. Superconductors are particularly well suited for THz detection because their characteristic energies (Josephson energy, superconducting gap, etc.) are naturally placed in the THz range and could be easily tuned for achieving optimal operation [13]. A large variety of superconducting THz detectors has been developed, including transition edge sensors [9], kinetic inductance detectors [7, 11] and hot-electron bolometers/mixers [8, 10].
@@
A current-biased Josephson junction (JJ) can act as a sensitive switching current detector (SCD) with a potential for single-photon resolution in a broad microwave (MW) to THz range [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13]. When JJ is biased just below the critical current, $I\lesssim I_{c}$, a small-amplitude electromagnetic signal triggers its switching from the superconducting to the resistive state. The responsivity of SCD can be very high, limited only by quantum fluctuations [13, 29]. The upper frequency depends on the characteristic voltage, $V_{c}=I_{c}R_{n}$, and ranges from sub-THz for low-$T_{c}$ [17], to THz for high-$T_{c}$ [26, 30, 27, 24] JJs. The highest $V_{c}\gtrsim 30\ \mathrm{mV}$ [31, 32] is achieved in intrinsic Josephson junctions (IJJs), naturally occurring in Bi2Sr2CaCu2O8+δ (Bi-2212) cuprates [33]. Operation of IJJs above 10 THz has been demonstrated [24, 30]. The atomic-scale of IJJs enables strong mutual coupling, advantageous for the creation of coherent THz electronics [34, 35, 36, 37, 38, 39, 40, 24, 27].
@@
Cascade multiplication is widely employed for enhancing detector sensitivity, e.g. in photo-multipliers and avalanche photodiodes [4]. For Josephson detectors, this could be achieved using arrays of mutually coupled JJs. The coupling leads to a current-locking phenomenon [41], where the switching of one JJ drags several neighbors into the resistive state, thereby multiplying the readout voltage [43, 41, 42]. A good detector should also have a high absorption efficiency, necessitating impedance matching with free space [46, 44, 45, 40].
@@
In this work we demonstrate novel types of cascade-amplified SCDs featuring arrays of coupled JJs. Collective, avalanche-like switching of many JJs leads to cascade amplification of the readout voltage, boosting the SCD sensitivity. We present two prototypes based on either linear arrays of low-$T_{c}$ Nb/NbxSi1-x/Nb JJs or stacked high-$T_{c}$ IJJs in Bi-2212 whiskers. We investigate both MW and THz responses and observe non-trivial MW polarization-loss diagrams, manifesting antenna effects associated with the sample geometry. This suggests good impedance matching and absorption efficiencies for both types of devices. We argue that the combination of high-$T_{c}$ superconductivity with high absorption efficiency and large cascade gain can yield broadband THz sensors with outstanding sensitivities.
@@ Results
JJ arrays can be advantageously used for photon detection. It has been shown that arrays can help in impedance matching and broaden the detector dynamic range [47, 48]. However, earlier studies were focusing on heterodyne mixers, operating at the quasiparticle branch of the current-voltage ($I$-$V$) characteristics. In contrast, here we consider SCD, where the signal is generated upon switching out of the superconducting state [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13].
Figure 1: Cascade SCD based on a linear Nb/NbxSi1-x/Nb array. (a) SEM image (false color) and layout of the array. The studied line with $N=128$ JJs is marked in red. (b) Ensemble of the $I$-$V$’s without MW irradiation, obtained at slightly different bias currents. The red oval indicates a collective self-resonant step. (c) The $I$-$V$’s with a fixed bias amplitude for different MW attenuations, $\mu$. (d) The voltage response (blue symbols, left axis) and the number of active JJs, $n$, (red line, right axis) as a function of MW power. (e) $V$ and $n$ as a function of the polarization angle, $\Theta$, for $P=\text{const}$. (f) Polarization loss-function. The orange line marks the array direction. Profound off-axis lobes manifest the traveling-wave antenna effect.
Our goal is to achieve cascade amplification of the readout voltage. Individual JJs inherently exhibit variations. Without interjunction coupling, the array $I$-$V$ is simply the sum of $I$-$V$’s of independent JJ. However, when the coupling is present, a current-locking phenomenon may occur [41], causing all (or several) JJs to switch simultaneously. Such collective switching leads to a cascade multiplication of the readout voltage, proportional to the number of current-locked JJs [43, 42, 41]. The efficiency of cascading depends on the uniformity of JJs and the strength of coupling. Therefore, arrays with many strongly interacting JJs are required.
@@
A linearly polarized MW signal at $f\simeq 74.5\ \mathrm{GHz}$ is delivered to the sample in a quasi-optical manner. MW attenuation, $\mu$, and polarization angle, $\Theta$, are adjusted using grid polarizers. The incoming MW power, $P$, is monitored by a Golay cell. Experimental details can be found in the Supplementary. All presented measurements are performed at zero magnetic field and $T\simeq$ 3.3 K.
@@
Fig. 1 (a) shows a scanning electron microscope (SEM) image and a layout of a $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ array. We study a linear section (marked red) with $N=128$ JJs. Details about fabrication and characterization of the arrays can be found in Refs. [49, 50, 51, 52]. Fig. 1 (b) shows the $I$-$V$s measured upon minor variation of the bias current amplitude, $I_{b}$, in the absence of MW. Multiple branches correspond to a different number, $n$, of active JJs. Interjunction coupling in these arrays is mediated by surface plasmons [51]. Such traveling waves are manifested by the appearance of collective resonant steps in the $I$-$V$’s [52]. The largest step is marked in the black $I$-$V$.
Fig. 1 (c) shows a set of the $I$-$V$’s measured at a fixed $I_{b}$ for different MW powers, characterized by the attenuation factor, $\mu$. With increasing $P$, the switching current, $I_{s}$, decreases and progressively more JJs switch into the resistive state, leading to increasing of the readout voltage. Blue symbols and the red line in Fig. 1 (d) represent the voltage response of the array and the number of active JJs, $n$, respectively. It can be seen that the rapid increase of $V$ at low $P$ is proportional to $n$, illustrating the cascade amplification phenomenon. At higher power all $N=128$ JJs switch to the resistive state and the responsivity $dV/dP$ is greatly reduced. In this case the response is solely due to the decrease of $I_{s}(P)$, qualitatively similar to a single-JJ SCD [24, 13]. Thus the large difference in low and high- power responsivities highlights the advantage of array-based cascade SCD compared to a single-JJ SCD.
@@
Figure 2: Microwave characteristics of a Bi-2212 mesa. (a) A SEM image of a whisker-based device. (b) Sketch of a mesa, forming a turnstile antenna. (c) Two $I$-$V$s of a mesa with $N=250$ IJJs, measured with different bias amplitudes without MW. (d) Ensemble of the $I$-$V$s for different MW powers, determined by the attenuation factor $\mu$, at constant $I_{b}=38.1~{}\mu\mathrm{A}$. (e) The mesa voltage (symbols) and the number of active IJJs, $n$, (lines) versus the MW power, for $I_{b}=25.3$ and $38.1~{}\upmu\mathrm{A}$. (f) The polarization-loss diagram. The blue line indicates the orientation of the whisker. The yellow background shows the large-scale electrode geometry. The four-fold symmetry is consistent with the turnstile-antenna geometry of the device.
@@
A sensitive detector should efficiently absorb the incoming power. The optimal absorption efficiency ($50\ \%$) is achieved at the impedance matching condition [45, 46, 44]. A JJ alone cannot effectively absorb radiation because its size is much smaller than the wavelength of radiation in free space, $\lambda_{0}$, [44]. Therefore, mm-long electrodes play a crucial role as impedance matching antennas [40]. The presence or absence of geometry-specific antenna effects provides a clue about the effectiveness of impedance matching and MW absorption.
@@
The active part of Nb-array is about $2\ \mathrm{mm}$ long ($\sim\lambda_{0}/2$ at 74.5 GHz). There are also passive electrodes (without JJ) with the total length $\sim 1\ \mathrm{cm}$. Long electrodes may act as non-resonant traveling wave antennas, which are characterized by asymmetric off-axis directivities [50, 51]. Fig.1 (e) shows the detector voltage (blue symbols) and the number of active JJs (red line) as a function of the MW polarization angle, $\Theta$, at a constant $P$. $\Theta=0$ corresponds to the MW electric field parallel to the array line. Fig. 1 (f) displays the corresponding polarization-loss diagram [53], $\Delta V/V_{G}(\Theta)$, where $\Delta V$ is the array voltage with respect to the turned-off MW, and $V_{G}\propto P$ is the Golay cell voltage (constant in this experiment). The diagram exhibits two lobes with profound maxima at $\Theta\simeq 50^{\circ}$ and $-130^{\circ}$ with respect to the array. The off-axis behavior is consistent with traveling-wave antenna operation of long electrodes, enabling good impedance matching and absorption efficiency.
@@
Figure 3: Detection of THz radiation from a Bi-2212 mesa. (a) The $I$-$V$ of the generator mesa for upward (blue) and downward (magenta) bias sweeps. Top axis indicates the anticipated Josephson frequency. (b) and (c) Low-bias $I$-$V$’s of the detector mesa with (c) showing the $I$-$V$ of the weak surface junction. (d) Measured detector responses versus $P_{gen}=I_{gen}V_{gen}$. The olive line represents a single surface IJJ response (multiplied by a factor 50). The magenta curve represents the cascade SCD response of the same mesa. (e) A set of detector $I$-$V$’s with $I_{b}=370~{}\mu$A at different bias points of the generator, marked by the same-color circles in (a) and (d).
@@
Fig. 2 (c) displays two $I$-$V$’s of a mesa ($\sim 5\times 5$ $\upmu\mathrm{m}^{2}$), containing $N\simeq 250$ IJJs, without MW. The blue curve, with $n=117$ active IJJs, is measured with $I_{b}$ slightly above the mean switching current, $I_{s}\sim 50~{}\upmu\mathrm{A}$. A multi-branch structure ends at $V\simeq 5$ V when all $N\simeq 250$ IJJ’s switch into the resistive state (red curve). The large readout voltage (5 V) is caused by the large $V_{c}\simeq 20\ \mathrm{mV}$ per IJJ. Switching of the first and the last IJJ occurs within an interval, $\sigma_{n}\sim 5\ \upmu\mathrm{A}$   $\sim 0.1~{}I_{s}$, representing the width of the switching current distribution [55, 13]. By varying $I_{b}$ in this narrow range, we can continuously tune $n$ from 1 to 250. The SCD response [13],
$\delta V\simeq-\frac{NV_{c}}{\sigma_{n}}\frac{\partial I_{s}}{\partial I_{MW}}\delta I_{MW},$ (1)
where $I_{MW}$ is the MW current induced by the impacting radiation. The response is enhanced by the large cascade gain, $N$, and narrow width, $\sigma_{n}$.
@@
Fig. 2 (d) shows the $I$-$V$s of another mesa at different MW powers and $I_{b}=38.1~{}\mu$A. The cascade gain depends on $I_{b}$. Fig. 2 (e) represents the MW power dependencies of $V$ (symbols) and $n$ (lines), measured at two biases. At lower $I_{b}=25.3\ \upmu\mathrm{A}$, the sensitivity, $dV/dP$, remains small up to some threshold, $P\sim 0.2~{}P_{max}$. At higher $I_{b}=38.1~{}\upmu\mathrm{A}$, there is no threshold and the sensitivity is high at $P=0$. With further increase of $I_{b}$, the sensitivity remains high at $P=0$ but saturates at successively smaller $P$, thus reducing the dynamic range [13]. Therefore, there is an optimal bias ($I_{b}=38.1\ \upmu\mathrm{A}$ in this case) with both high sensitivity and large dynamic range. For comparison, in the Supplementary we analyze the ordinary SCD operation on the same mesa using a single surface IJJ. The surface IJJ has much smaller $I_{s}$ [32] and can be measured without activation of other IJJs [24]. The maximum single-junction response is almost 1000 times smaller than for the cascade SCD in Fig. 2 (e) at the same MW power for the same mesa.
@@
Fig. 2 (f) shows the polarization-loss diagram of this mesa. It has a four-fold shape with two smaller lobes aligned with the whisker (indicated by the blue line) and the larger lobes approximately perpendicular, slightly inclined in the direction of one of bias electrodes (see the Supplementary for more details). The large-scale electrode geometry of this device is shown in the background (yellow). The observed four-fold diagram is consistent with the turnstile-antenna geometry of the device, as shown in Fig. 2 (b) [56].
@@
To verify THz operation, we perform an in-situ generation-detection experiment in which one mesa emits Josephson radiation and a nearby mesa detects it [24, 27]. Fig. 3 (a) shows the $I$-$V$ of a generator mesa ($\sim 10\times 15~{}\mu$m2), containing $N=160\pm 10$ IJJs. We analyze the downturn part of it (magenta) at which all $N$ IJJs are active. In this case, the Josephson frequency is well defined, $f_{J}=2eV/hN$, as indicated in the top axis. The emission occurs at cavity modes in the mesa [24, 37, 38]. The key signature of such emission (as opposed to heating) is the non-monotonous dependence of the signal on the dissipation power, $P_{\text{gen}}=I_{\text{gen}}V_{\text{gen}}$, with distinct peaks at cavity resonances [24].
Fig. 3 (b) shows a low-bias part of the $I$-$V$ of a detector mesa ($\sim 15\times 15~{}\mu$m2), about $22~{}\mu$m away from the generator mesa. Fig. 3 (c) is a close-up on the central part, representing the $I$-$V$ of the surface junction. It has $I_{c}^{\prime}\simeq 15~{}\mu$A, significantly smaller than $I_{c}\simeq 370~{}\mu$A for the rest of the IJJs. This disparity allows ordinary SCD measurements at $I_{c}^{\prime}<I_{b}\ll I_{c}$, which activates only one surface IJJ [24]. The cascade SCD can be analyzed on the same mesa by increasing bias to $I_{b}\sim I_{c}$.
Fig. 3 (d) shows single-junction (olive, ×$\times 50$) and cascade (magenta) SCD responses versus $P_{gen}$, measured at bias amplitudes, $I_{b}=25$ and 370 $\mu$A, respectively. It can be seen that (in this case) the resolution of a single-junction SCD is not sufficient for resolving the emission. The cascade SCD demonstrates a much larger response with clearly visible emission peaks. The main emission for this generator mesa occurs in the range from $\sim 4$ to 8.4 THz, as indicated in the top axis. Fig. 3 (e) shows the detector $I$-$V$’s measured with $I_{b}=370~{}\mu$A at several bias points of the generator, marked by the same-color circles in Figs. 3 (a) and (d). It can be seen that cascade amplification factor, $n$, is playing the dominating role in the detector response.

### v2

@@ Abstract
Cascade multiplication is a common technique to enhance the sensitivity of photon detectors. In this study, we demonstrate novel cascade-amplified superconducting detectors utilizing arrays of Josephson junctions. The mutual coupling between junctions induces avalanche-like switching of multiple junctions upon photon absorption, leading to cascade amplification of the readout voltage. We present two prototypes featuring either low-$T_{c}$ linear $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Both devices exhibit clear antenna effects in microwave directivities, indicating good impedance matching and absorption efficiency. The combination of high absorption efficiency and large cascade amplification has the potential to produce broadband THz sensors with sensitivity exceeding $10^{13}\ \mathrm{V/W}$.
@@ Introduction
THz detectors find diverse applications, including security, environmental monitoring, medical imaging, chemical analysis, and fundamental research [1, 2, 3, 4]. Cooling to cryogenic temperatures improves detector sensitivity by reducing thermal noise and background radiation. Ultra-sensitive cryogenic detectors based on semiconductors [5], graphene [6], and superconductors [7] have been demonstrated. Superconductors are particularly well suited for THz detection because their characteristic energies (Josephson energy, superconducting gap, etc.) are naturally placed in the THz range and could be easily tuned for achieving optimal operation [8].
@@
A Josephson junction (JJ) can act as a sensitive switching current detector (SCD) with a potential for single-photon resolution in a broad microwave (MW) to THz range [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 8]. When JJ is biased just below the critical current, $I\lesssim I_{c}$, a small-amplitude electromagnetic signal triggers its switching from the superconducting to the resistive state. The responsivity of SCD can be very high, limited only by quantum fluctuations [8, 24]. The upper frequency depends on the characteristic voltage, $V_{c}=I_{c}R_{n}$, and ranges from sub-THz for low-$T_{c}$ [12], to THz for high-$T_{c}$ [21, 25, 22, 19] JJs. The highest $V_{c}\gtrsim 30\ \mathrm{mV}$ [26, 27] is achieved in intrinsic Josephson junctions (IJJs), naturally occurring in Bi2Sr2CaCu2O8+δ (Bi-2212) cuprates [28]. Operation of IJJs above 10 THz has been demonstrated [19, 25]. The atomic-scale of IJJs enables strong mutual coupling, advantageous for the creation of coherent THz sources [29, 30, 31, 32, 33, 34, 35, 19, 22].
@@
Cascade multiplication is widely employed for enhancing detector sensitivity, e.g. in photo-multipliers and avalanche photodiodes [4]. For Josephson detectors, this could be achieved using arrays of mutually coupled JJs. The coupling leads to a current-locking phenomenon [36], where the switching of one JJ drags several neighbors into the resistive state, thereby multiplying the readout voltage [38, 36, 37]. A good detector should also have a high absorption efficiency, defined as the ratio of absorbed to incoming THz power. This necessitates impedance matching with free space, which is equally important for both THz sources and sensors [41, 39, 40, 35].
@@
In this work we demonstrate novel types of cascade-amplified SCDs featuring arrays of coupled JJs. Collective, avalanche-like switching of many JJs leads to cascade amplification of the readout voltage, boosting the SCD sensitivity. We present two prototypes based on either linear arrays of low-$T_{c}$ Nb/NbxSi1-x/Nb JJs or stacked high-$T_{c}$ IJJs in Bi-2212 whisker-type crystals. We investigate both MW and THz responses and observe non-trivial MW directivities, manifesting antenna effects associated with the sample geometry. This suggests good impedance matching and absorption efficiencies for both types of devices. We argue that the combination of high-$T_{c}$ superconductivity with high absorption efficiency and large cascade gain can yield broadband THz sensors with outstanding sensitivities.
@@ Results
JJ arrays can be advantageously used for photon detection. It has been shown that utilization of arrays can help in impedance matching and broaden the detector range [42, 43]. However, earlier studies were focusing on heterodyne mixers, operating at the quasiparticle branch of the current-voltage ($I$-$V$) characteristics. In contrast, here we consider SCD, where the signal is generated upon switching out of the superconducting state [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 8].
Figure 1: Cascade SCD based on a linear Nb/NbxSi1-x/Nb array. (a) SEM image (left, false color) and layout (right) of the array. The studied line is marked in red. (b) Ensemble of the $I$-$V$’s without MW irradiation, obtained at slightly different bias currents. (c) The $I$-$V$’s with a fixed bias amplitude for different MW powers at $f\simeq 74.5\ \mathrm{GHz}$. Switching of multiple JJs leads to cascade amplification of the readout voltage. (d) The lock-in resistance of the array, $R$, (blue symbols, left axis) and the number of active JJs, $n$, (red line, right axis) as a function of MW power. (e) $R$ and $n$ as a function of the polarization angle, $\Theta$, for $P=\text{const}$. (f) Angular diagram of detector sensitivity. The orange line marks the orientation of the array. A profound off-axis directivity manifest the traveling-wave antenna effect.
Our goal is to achieve cascade amplification of the readout voltage. Individual JJs inherently exhibit variations. Without interjunction coupling, the array $I$-$V$ is simply the sum of independent JJ $I$-$V$’s. However, when the coupling is present, a current-locking phenomenon may occur [36], causing all (or several) JJs to switch simultaneously. Such collective switching leads to a cascade multiplication of the readout voltage, proportional to the number of current-locked JJs [38, 37, 36]. The efficiency of cascading depends on the uniformity of JJs and the strength of coupling. Therefore, arrays with many strongly interacting JJs are required to achieve high gain.
@@
A linearly polarized MW signal at $f\simeq 74.5\ \mathrm{GHz}$ is delivered to the sample in a quasi-optical manner. MW attenuation and polarization angle are adjusted using grid polarizers. The incoming MW power, $P$, is monitored by a Golay cell. Experimental details can be found in the Supplementary. All presented measurements are performed at zero magnetic field and $T\simeq$ 3.3 K.
@@
Fig. 1 (a) shows a scanning electron microscope (SEM) image and a layout of a $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ array. We study a linear section (marked red) with $N=128$ JJs. Details about fabrication and characterization of the arrays can be found in Refs. [44, 45, 46, 47]. Fig. 1 (b) shows the $I$-$V$s measured upon minor variation of the bias current amplitude, $I_{b}$, in the absence of MW. Multiple branches correspond to a different number, $n$, of active JJs. Interjunction coupling in these arrays is mediated by surface plasmons [46]. Such traveling waves are manifested by the appearance of collective resonant steps in the $I$-$V$’s [47]. The largest step is clearly visible in the rightmost (black) $I$-$V$ in Fig. 1 (b) with all $N=128$ active JJs.
@@
Figure 2: Microwave characteristics of a Bi-2212 mesa. (a) A SEM image of a whisker-based device. (b) Sketch of a mesa formed at the intersections of the whisker with a top Au-electrode. It has a turnstile antenna geometry. (c) Two $I$-$V$s of a mesa with $N=250$ IJJs, in the absence of MW. (d) Ensemble of the $I$-$V$s for different MW powers at $I_{b}=38.1~{}\mu\mathrm{A}$. (e) The mesa resistance (symbols) and the number of active IJJs, $n$, (lines) versus the MW power, for $I_{b}=25.3$ and $38.1~{}\upmu\mathrm{A}$. (f) The directivity diagram. The blue line indicates the orientation of the whisker. The yellow background shows the large-scale electrode geometry. The four-fold directivity is consistent with the turnstile-antenna geometry of the device.
Fig. 1 (c) shows a set of the $I$-$V$’s measured at a fixed $I_{b}$ for different MW powers. With increasing $P$, the switching current, $I_{s}$, decreases and progressively more JJs switch into the resistive state, leading to increasing of the readout voltage. Blue symbols and the red line in Fig. 1 (d) represent the lock-in resistance of the array, $R$, and the number of active JJs, $n$, respectively. It can be seen that the rapid increase of $R$ at low $P$ is proportional to $n$, illustrating the cascade amplification phenomenon. At high power, when all $N=128$ JJs are in the resistive state, the responsivity $dR/dP$ is greatly reduced. In this case the response is solely due to the decrease of $I_{s}(P)$, qualitatively similar to a single-JJ SCD [19, 8]. Thus the large difference in low and high- power responsivities highlights the advantage of array-based cascade SCD compared to a conventional single-JJ SCD.
@@
A sensitive detector should efficiently absorb the incoming power. The optimal absorption (and emission) efficiency ($50\ \%$) is achieved at the impedance matching condition [40, 41, 39]. A JJ alone cannot effectively absorb radiation because its size is much smaller than the wavelength of radiation in free space, $\lambda_{0}$, [39]. Therefore, mm-long electrodes play a crucial role as impedance matching antennas [35]. The presence or absence of geometry-specific antenna effects provides a clue about the effectiveness of impedance matching and MW absorption.
@@
The active part of Nb-array is about $2\ \mathrm{mm}$ long ($\sim\lambda_{0}/2$ at 74.5 GHz). Additionally, there are long passive electrodes (without JJ) with a total length $\sim 1\ \mathrm{cm}$. These long electrodes may act as non-resonant traveling wave antennas, which are characterized by asymmetric of-axis directivities [45, 46]. Fig.1 (e) shows the array resistance (blue symbols) and the number of active JJs (red line) as a function of the polarization angle, $\Theta$, at a constant $P$. $\Theta=0$ corresponds to the MW electric field parallel to the array line. Fig. 1 (f) displays the corresponding sensitivity diagram, $\Delta R/V_{G}(\Theta)$, where $\Delta R$ is the resistance increase with respect to the turned-off MW, and $V_{G}\propto P$ is the Golay cell voltage (constant in this experiment). The diagram exhibits two lobes with a large directivity at $\Theta\simeq 50^{\circ}$ with respect to the array. Such off-axis directivity indicates traveling-wave antenna operation of long electrodes, enabling good impedance matching and absorption efficiency.
Figure 3: Detection of THz radiation from a Bi-2212 mesa. (a) The $I$-$V$ of the generator mesa for upward (blue) and downward (magenta) bias sweeps. Top axis indicates the anticipated Josephson frequency. (b) and (c) Low-bias $I$-$V$’s of the detector mesa with (c) representing the $I$-$V$ of the weak surface junction. (d) Measured detector responses versus $P_{gen}=I_{gen}V_{gen}$. The olive line is obtained with $I_{b}=25~{}\mu$A, activating only a single surface IJJ. The magenta curve is measured with $I_{b}=370~{}\mu$A, activating many IJJs. (e) A set of detector $I$-$V$’s with $I_{b}=370~{}\mu$A at different bias points of the generator, marked by the same-color circles in (a) and (d). It can be seen that the cascade amplification factor, $n$, is playing the dominant role in boosting the detector sensitivity.
@@
Fig. 2 (c) displays two $I$-$V$’s of a mesa ($\sim 5\times 5$ $\upmu\mathrm{m}^{2}$), containing $N\simeq 250$ IJJs, without MW. The blue curve, with $n=117$ active IJJs, is measured with $I_{b}$ slightly above the mean switching current, $I_{s}^{*}\sim 50~{}\upmu\mathrm{A}$. A multi-branch structure ends at $V\simeq 5$ V when all $N\simeq 250$ IJJ’s switch into the resistive state (red curve). The large readout voltage is caused by the large $V_{c}\simeq 20\ \mathrm{mV}$ per IJJ. Switching of the first and the last IJJ occurs within an interval, $\sigma\sim 5\ \upmu\mathrm{A}$   $\sim 0.1~{}I_{s}^{*}$, representing the width of the switching current distribution [49, 8]. By varying $I_{b}$ in this narrow range, we can continuously tune $n$ from 1 to 250. The responsivity of SCD is determined by the effective differential resistance, $R_{d}=dV/dI$, in the switching region [9]. For this mesa it is very large, $R_{d}=NV_{c}/\sigma\sim 1\ \text{M}\Omega$, as a consequence of the cascading.
@@
Fig. 2 (d) shows the $I$-$V$s of another mesa at different MW powers and $I_{b}=38.1~{}\mu$A. The cascade gain depends on bias. Fig. 2 (e) represents the MW power dependencies of $R$ (symbols) and $n$ (lines), measured at two biases. At lower $I_{b}=25.3\ \upmu\mathrm{A}$, the sensitivity, $dR/dP$, remains small up to some threshold, $P\sim 0.2~{}P_{max}$. At higher $I_{b}=38.1~{}\upmu\mathrm{A}$, there is no threshold and the sensitivity is high at $P=0$. With further increase of $I_{b}$, the sensitivity remains high at $P=0$ but saturates at successively smaller $P$, thus reducing the dynamic range [8]. Therefore, there is an optimal bias ($I_{b}=38.1\ \upmu\mathrm{A}$ in this case) with both high sensitivity and large dynamic range. For comparison, in the Supplementary we analyze the ordinary SCD operation on the same mesa using a single surface IJJ. The surface IJJ has much smaller $I_{s}$ [27] and can be measured without activation of other IJJs [19]. The maximum single-junction response is $\Delta R\sim 40~{}\Omega$, much smaller than $\sim 8~{}\mathrm{k}\Omega$ for the cascade SCD in Fig. 2 (e) at the same MW power for the same mesa.
@@
Fig. 2 (f) shows the MW directivity of this mesa. It exhibits a four-fold shape. The two smaller lobes are aligned with the whisker (indicated by the blue line). The larger lobes are approximately perpendicular, slightly inclined in the direction of one of bias electrodes (see the Supplementary for more details). The large-scale electrode geometry of this device is shown in the background (yellow). The observed four-fold directivity is consistent with the whisker-mesa-electrode geometry, which form a turnstile-antenna, as shown in Fig. 2 (b) [50].
@@
To verify THz operation, we perform an in-situ generation-detection experiment in which one Bi-2212 mesa emits Josephson radiation and a nearby mesa detects it [19, 22]. Fig. 3 (a) shows the $I$-$V$ of a generator mesa ($\sim 10\times 15~{}\mu$m2), containing $N=160\pm 10$ IJJs. We analyze the downturn part of it (magenta) at which all $N$ IJJs are active. In this case, the Josephson frequency is well defined, $f_{J}=2eV/hN$, as indicated in the top axis. Emission from Bi-2212 mesas occurs at cavity modes inside the mesa [19, 32, 33]. The key signature of such emission (as opposed to heating) is the non-monotonous dependence on the dissipation power, $P_{\text{gen}}=I_{\text{gen}}V_{\text{gen}}$, with distinct peaks at cavity resonances [19].
Fig. 3 (b) shows a low-bias part of the $I$-$V$ of a detector mesa ($\sim 15\times 15~{}\mu$m2), about $22~{}\mu$m away from the generator mesa. Fig. 3 (c) is a close-up on the central part, representing the $I$-$V$ of the surface junction. It has $I_{c}^{\prime}\simeq 15~{}\mu$A, significantly smaller than $I_{c}\simeq 370~{}\mu$A for the rest of the IJJs. This disparity allows ordinary SCD measurements at $I_{c}^{\prime}<I_{b}\ll I_{c}$, which activates only one surface IJJ [19]. The cascade SCD can be analyzed on the same mesa by increasing bias to $I_{b}\gtrsim I_{c}$.
Fig. 3 (d) shows single-junction (olive) and cascade (magenta) SCD responses versus $P_{gen}$, measured at bias amplitudes, $I_{b}=25$ and 370 $\mu$A, respectively. It can be seen that (in this case) the resolution of a single-junction SCD is not sufficient for resolving the emission. The cascade SCD demonstrates a much larger response with clearly visible emission peaks. The main emission for this generator mesa occurs in the range from $\sim 4$ to 8.4 THz, as indicated by the top axis in Fig. 3 (d). Fig. 3 (e) shows the detector $I$-$V$’s measured with $I_{b}=370~{}\mu$A at several bias points of the generator, marked by the same-color circles in Figs. 3 (a) and (d). It can be seen that cascade amplification factor, $n$, is playing the dominating role in the detector response.